Use Euler’s method to approximate the solution of the ODE
dx/dt = t − x,              x(0) = 1
up to time t = 0.5 with a step size of h = 0.1. Find the actual solution of the equation and graph the approximate solution and the actual solution.
Answer #1

Given thatー-11, x)-t-r with initial condition x(0) = 1 i.e., = 0, Take h =0.1 The Eulers method is石+,-石+hf (4 ,%) =1 而=xo +k